# Dynamical Logical Qubits in the Bacon-Shor Code

### v1

@@ IV.1 Decoding and Code Distance
For simplicity, we will assume the case of a single dynamical logical qubit, as the extension to multiple logical qubits is straightforward, and briefly discussed towards the end of this section. In addition, note that each of the ISGs is equivalent to a clockwise $\pi/2$ rotation of the previous ISG, together with interchanging $X\leftrightarrow Z$, so that it suffices to discuss errors occuring between any two consecutive rounds.
Next, suppose concretely that we are in the ISG $S^{(4k)}$ ($k\geq 1$), and suppose that a single qubit $X$ error occurs on some row in the bulk, far from the top and bottom boundaries of the square lattice, and far from the boundaries of the cells A, B, C and D. This lone $X$ error will flip the two permanent $Z$-type stabilizers that straddle this row, and that are measured in the next row. If this $X$ error occurred on the right of the blue domain wall along plaquette column AD, then it will also additionally flip the two transient stabilizers that exist along this defect column, and that also straddle this row and are measured in the next round. Thus, such errors give rise to hyper-edges connecting 4 vertices in the decoding hyper-graph.
At the top and bottom boundary rows, we may add additional imaginary stabilizers that get flipped in addition to the lone stabilizers, permanent and transient, in order to apply matching algorithms. This leaves us with three rows left to consider: the top row of cells AB, the row between cells AB and CD, and the bottom row of cells CD. Note that an $X$ error on the row between cells AB and CD is ISG equivalent to a single $X$ error anywhere on that row, so that we may treat it as occuring e.g. on the left-most column.
Between rounds $4k$ and $14k+$ 1, an $X$ error on the bottom row of cells CD can be identified in a somewhat similar fashion, as it flips the pairs of permanent $Z$-type stabilizers that straddle this row and that all get measured in the round $14k+$ 1. In addition, it also flips the transient $Z$-type stabilizer associated with the cell below cell D, which also gets measured in the next round $14k+$ 1. Although the $\overline{Z}$ operator of cell D gets measured in round $14k+$ 1, it cannot be used as a syndrome bit as it has not persisted through the previous round, nor does it persist in the subsequent round. In general, none of the $\overline{Z}$ operators of cells/gauge qubits A, B, C or D may be used as syndrome bits, even if they exist in some of the ISGs. We therefore add imaginary boundary transient $Z$-type stabilizers at cells A, B, C and D to allow for matching with neighboring transient $Z$-type stabilizers.
To handle the remaining two cases, we may adopt two different decoding strategies, one that detects and fixes errors as they accumulate, i.e. in the same time slice, and the second where non-trivial syndrome measurements, even with perfect measurements, are edge-connected across time slices separated by two measurement rounds. Using the first strategy, a single $X$ error on either one of the above two rows, i.e. the row above cells AB, or the row between cells AB and CD, will flip a single permanent $Z$-type stabilizer that borders the plaquette row AB in the next round $14k+$ 1. In addition, an $X$ error on the row above cells AB will additionally flip a single transient $Z$-type stabilizer above cell A. In the first decoding strategy, we associate an imaginary permanent $Z$-type stabilizer to the plaquette row AB, as well as an imaginary transient $Z$-type stabilizer to the cell A, to enable pairing with the nearby stabilizers in a matching algorithm. This effectively reduces the distance by a factor of roughly 1/2, as an error chain of weight $\lesssim d/2$ consisting of $X$ errors on all the rows above the top row of cells AB will produce the same syndrome information as a single $X$ error on the row above cells AB.
Such a chain would however be distinguished from the single $X$ error if one also includes the $Z$-type stabilizer along plaquette row AB, but this does not get measured until measurement round $34k+$ 3. In the second decoding strategy, the decoding graph consists of edges that connect the two permanent $Z$-type stabilizers that border plaquette row AB, measured at round $14k+$ 1, to the permanent $Z$-type stabilizer that runs along plaquette row AB. Together with the flipped transient $Z$-type stabilizer above cell B, as well as the flipped imaginary transient stabilizer at cell B, this enables distinguishing the lone $X$ error from the error chain described above, and therefore restores the code distance to $\sim d$. In either case, the code distance remains $\Theta(d)$.
@@
We note that there currently exists no general well-formed notion of the code distance of a Floquet code. In particular, it is not simply the minimum code distance of any of the ISGs considered as stabilizer codes in isolation, as the time dynamics may allow for small weight errors to evolve into logical errors, even if the ISGs have large code distances by themselves. For one thing, we can multiply instantaenous logical operators by not only instantaneous stabilizer elements, but also elements of the ISGs before and after the current round. This is so because an error of the form $\mathcal{L}^{(r)}s^{(r-1)}$ between measurement rounds $r-1$ and $r$, where $\mathcal{L}^{(r)}$ is a logical operator at round $r$, and $s^{(r-1)}\in\mathcal{S}^{(r-1)}$, has the same effect as an error of the form $\mathcal{L}^{(r)}$ at round $r$. Similarly, an error of the form $s^{(r+1)}\mathcal{L}^{(r)}$ between measurement rounds $r$ and $r+1$ will also have the same effect. In the present case, such transformations do not affect the code distance scaling reported here.
@@
Even more generally speaking, we may say that two Pauli operators $\sigma_{\alpha}$ and $\sigma_{\beta}$ are space-time connected in a Floquet code if their difference is a a chain of instantaneous stabilizer elements, i.e. $\sigma_{\alpha}=s_{r+t}\dots s_{r}\sigma_{\beta}$. It may be the case that a low weight Pauli error $\sigma_{\beta}$ traverses, in its spacetime path, an undetectable logical error, before evoling into a partially detectable error, whose correction does not fix the previous logical error, or even getting absorbed into the ISG. Such a logical error occurs in [6], for instance. It may therefore be desirable to keep the number of time slices between which stabilizers are measured and used as syndromes to be as minimal as possible when desigining Floquet codes, which motivates the distance $\sim d/2$ decoding strategy we described above. In the absence of a formal definition of the code distance of a Floquet code, we simply identify the code distance as the smallest weight of any error operator that does not produce any non-trivial syndromes, given the decoding hyper-graph, which in turn depends on the choice of the specific error model we consider.
@@
We also note that like the parent Bacon-Shor subsystem code, the Floquet-Bacon-Shor family of codes introduced here do not possess a threshold. This is attributable to the fact that the weight of the stabilizers used in both examples of codes grows as the distance increases, so that erroneous stabilizer measurements get likelier with increasing distance for any fixed value of independent error probability.
@@ IV.2 Self-correcting errors
Suppose that a 2-qubit $XX$ error occurs on an edge somewhere above cell A between rounds 0 and 1 mod 4. This will flip the two transient, but not the two permanent, $Z$-type stabilizers that straddle this edge, lie along column AD, and are measured in round 1 mod 4. In the very next round 2 mod 4, we carry out a measurement of an $XX$ check operator on precisely the same edge, and the error becomes part of the ISG. In other words, the appropriate correction operator automatically applies itself because of the measurement schedule itself. This occurs even if this 2-qubit error anti-commutes with a logical operator at round 0 mod 4, such as if it occurred on the top edge of cell A. We show this more explicitly in the appendix.
@@ IV.3 Multiple Dynamical Qubits
The discussion above generalizes fairly straightforwardly to the case of adding multiple dynamical logical qubits. In this case, the main difference is that the hyper-edges in the decoding hyper-graph now connect more vertices. For instance, if we have 4 dynamical logical qubits arranged in a manner such that the pairs of cells $(A_{0},B_{0})$ and $(A_{1},B_{1})$ exist in the same plaquette row, and the pairs $(D_{0},C_{0})$ and $(D_{1},C_{1})$ exist in the plaquette row below it, then a single qubit $X$ error on some row in the bulk towards the right of defect column $A_{1}D_{1}$ between rounds 0 and 1 mod 4 will flip the values of the two permanent $Z$-type stabilizers, the two transient $Z$-type stabilizers along column $A_{1}D_{1}$, as well as the two transient $Z$-type stabilizers along column $A_{0}D_{0}$, that straddle this row. Thus, such an error gives rise to a hyper-edge connecting 6 vertices in the decoding hyper-graph.
The only other major difference is that an even number of $X$ errors between gauge defects in the same plaquette row are now part of the ISG at round 0, and no longer count as errors. Only single qubit $X$ errors between any consecutive pair of gauge defects, or a gauge defect and a lattice boundary count as errors, and similarly for $Z$ errors.
@@ V Conclusions
In this paper, we have described the introduction of dynamical logical qubits to the Bacon-Shor subsystem code by modifying the measurement schedule of the weight 2 check operators that generate the gauge group of the parent sub-system code. Such a measurement schedule introduces and maintains gauge defects throughout the measurement cycle. This broad prescription of identifying the virtual qubit degrees of freedom in a subsystem code, and picking a measurement schedule that can maintain gauge defects across measurement cycles as a way to introduce dynamical logical qubits may be more broadly applicable in constructing more examples of Floquet codes. We showed that independent errors of a fairly simple noise model naturally lead us to consider a decoding hyper-graph, which is generally more complicated to work with than decoding graphs that arise in surface codes, for instance. This additional complication in the decoding task is also shared to some degree with LDPC codes, in which independent errors, or faults, also typically flip more than two stabilizer syndromes, and progress in constructing good decoders for LDPC codes may also benefit decoding Floquet codes, such as the one describe here. We leave a detailed consideration of fault tolerance, together with fault tolerant implementation of logical gates, in this code to future work. Lastly, we also note that skipping measurements of some check operators in some rounds was also used in [30] to introduce a threshold to the Bacon-Shor code. In our paper, we used such skipped measurements to introduce dynamical logical qubits to the Bacon-Shor code. We leave it as an open problem to investigate whether one could combine both sets of ideas to introduce a threshold to the Bacon-Shor code in the presence of dynamical logical qubits.
@@ VI Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. M.S.A. acknowledges support from USRA NASA Academic Mission Services under contract No. NNA16BD14C with NASA, with this work funded under the NASA-DOE interagency agreement SAA2-403602 governing NASA’s work as part of the SQMS center. The authors wish to thank Jason Saied and Ryan LaRose for useful discussions on Floquet and error correcting codes in general, as well as Nikolas Breuckmann, Christophe Vuillot and Margarita Davydova for helpeful comments during the final stages of this work.

### v2

@@ IV.1 Decoding and Code Distance
For simplicity, we will consider the case of a single dynamical logical qubit; the the extension to multiple logical qubits is straightforward, as briefly discussed towards the end of this section. In addition, since each of the ISGs is equivalent to a clockwise $\pi/2$ rotation of the previous ISG, together with interchanging $X\leftrightarrow Z$, it suffices to discuss errors occuring between any two consecutive rounds.
We have a choice with respect to a decoding approach. We could form detectors out of the permanent Bacon-Shor stabilizers along with the gauge fixed operators along the defect rows/columns. This approach, however, in general leads to a decoding hyper-graph, which is typically more complicated to decode then decoding graphs. To see that this approach leads to a hypergraph, consider the ISG at round $4k$ and a single qubit $X$ error anywhere to the right of the defect column AD away from the gauge defect or the lattice boundary. Any such error would flip both the pair of permanent stabilizers and the gauge fixed operators that straddle the row along which this error occurred. Thus, such an error would lead to a hyper-edge in the decoding hyper-graph that connects all $4$ of these stabilizers.
To avoid creating a hypergraph, a useful strategy is to consider the transient stabilizers together with the product of permanent and transient stabilizers in the same plaquette row/column as the error syndromes. For instance, for the ISG $S^{(4k)}$ $(k\geq 1)$, for any square cell/plaquette along the plaquette column AD, except cells A and D, we would take the product of all $Z$ checks to the right of this cell as one stabilizer syndrome and the product of all $Z$ checks to the left of this cell as another stabilizer syndrome. Both such stabilizers are transient; they only survive until the next round. In this example, they serve as syndromes between round $14k-$ 1, when they are initialized, and $14k+$ 1, when they are re-measured, persisting through the intermediary round $4k$. Single-qubit $X$ errors on any row, except the top and bottom rows or those bordering either cell A or D, will flip a nearest-neighbor pair of such transient stabilizers. A single-qubit $X$ error on either the top or bottom most row will flip a single such transient stabilizer, depending on which side of the domain wall it occurred. We also have permanent $Z$-type stabilizers along plaquette rows AB and CD, which are measured in rounds $14k-$ 1 and $14k+$ 1 respectively.
Figure 5: (Top) ISG progression reproduced for convenience. Below, we depict the subgraph of the decoding graph that can detect errors occurring between rounds $14k+$ 1 and $14k-$ 1. All nodes represent $Z$ type stabilizers. Circular nodes depict temporary stabilizers that get initialized and measured at round $14k-$ 1, and then subsequently re-measured to form detectors and then destroyed at round $14k+$ 1. Square nodes represent the permanent Bacon-Shor stabilizers, which get measured once per measurement cycle. (a) Decoding using only the information available at the current time slice will necessarily exclude one of the permanent Bacon-Shor stabilizers, resulting in a disconnected decoding graph, and reducing the weight to $\sim d/2$, the length of either one of the straight line sub-graphs shown above. (b) By taking into account information syndrome information received 2 rounds later, one can form edges in the decoding graph between the permanent Bacon-Shor stabilizer measured along row AB at round $34k+$ 3 (colored in violet), and the rest of the stabilizers. Edges depicted above between the violet and blue nodes cut across time slices separated by 2 measurement rounds.
One decoding strategy is to detect and correct errors immediately, in the same time slide as they are detected. To see the limitation of this approach, consider round $14k+$ 1, when in addition to measuring the transient $Z$ stabilizers along every plaquette row except AB and CD, we also measure the permanent $Z$ stabilizer along row CD. This approach would result in a disconnected decoding graph with two components, one connecting the measured stablizers above the plaquette row AB and the other connecting the measured stabilizers below the plaquette row AB. In this approach, however, an error chain of weight roughly half the lattice length would produce a trivial syndrome, resulting in a code distance of only $\sim d/2$. Such a chain could be distinguished from the single $X$ error if one also includes the $Z$-type stabilizer along plaquette row AB, but this does not get measured until measurement round $34k+$ 3.
This observation suggests that a better decoding strategy would make use of information from multiple time slices. Observe that a single qubit $X$ error on any (vertex) row between the two plaquette rows AB and CD will flip the two permanent $Z$ stabilizers associated with these rows, separated by 2 measurement rounds. If we include such edges in the decoding graph, we can correct errors of weight $\sim d/2$, thereby restoring the code distance to $\sim d$. In any case, all single qubit errors lead to defects in either one or two stabilizers, and therefore standard matching techniques can be applied for the decoding task. The structure of the decoding graph is illustrated in Fig. 5.
The code distance we report here is the smallest weight of any non-trivial operator that produces a trivial syndrome, given the decoding graph, which in turn depends on the error model described above. Computing the distance of a code is generally NP hard, and we expect it remains so for Floquet codes. Moreoever, the notion of distance for a Floquet code is more nuanced than for a stabilizer or even subsystem codes. To begin with, the code distance is not simply the minimum code distance of any of the ISGs considered as stabilizer codes in isolation, as the time dynamics may allow for small weight errors to evolve into logical errors, even if the ISGs have large code distances by themselves.
@@
Additionally, we can multiply instantaneous logical operators not only by instantaneous stabilizer elements, but also elements of the ISGs before and after the current round. This is so because an error of the form $\mathcal{L}^{(r)}s^{(r-1)}$ between measurement rounds $r-1$ and $r$, where $\mathcal{L}^{(r)}$ is a logical operator at round $r$, and $s^{(r-1)}\in\mathcal{S}^{(r-1)}$, has the same effect as an error of the form $\mathcal{L}^{(r)}$ at round $r$. Similarly, an error of the form $s^{(r+1)}\mathcal{L}^{(r)}$ between measurement rounds $r$ and $r+1$ will also have the same effect. In the present case, such transformations do not affect the code distance scaling reported here. One way to see this is to note that the union of the measurements in two consecutive rounds of measurement forms a subsystem code. The minimum distance of all such subsystem codes across the entire measurement cycle provides a better upper bound to the distance of a Floquet code, than the minimum ISG distance, but in the general case it remains only an upper bound.
The above two upper bounds to the distance of a Floquet code are described more formally in [39] (which appeared a day after the initial posting of this paper on the arXiv preprint server), where they are referred to as the ISG and subsystem distances respectively. Additionally, [39] describes the notion of an unmasked distance, which provides an even better upper bound to the code distance. The unmasked distance depends on the notion of unmasked stabilizers, which are roughly speaking the stabilizers from any given ISG that survive a given measurement sequence, and are able to form detectors. In the case of the Floquet-Bacon-Shor codes we have described, the distance algorithm of [39] identifies precisely such stabilizers, and predicts the same code distance $\sim d$ that we argued for above using the decoding graph perspective.
In the case of the ISG at round(s) $4k$, these detectors are formed by the $\overline{X}$ ($\overline{Z}$) operators of all the gauge qubits along the plaquette row AB (column AD), except for gauge qubits A and B (A and D), as well as their products with the permanent Bacon-Shor stabilizers in the same row (column), along with the permanent Bacon-Shor stabilizers along column AD and row AB. Meanwhile, the gauge group consists of all other gauge fixed $\overline{X}$ operators at round $4k$ as well as the gauge $\overline{Z}$ operators that destabilize the former in the next round $14k+$ 1. We provide a detailed calculation using the distance algorithm of [39] applied to an example Floquet-Bacon-Shor code in Appendix C.
@@
As the authors state, the distance algorithm of [39] only works in the case of perfect measurements, so to restore our code distance to $O(d)$ in the presence of measurement errors, we would need to repeat each measurement round $O(d)$ many times, as argued above. Moreover, even without measurement errors, the algorithm may not be able to capture the propagation of low-weight errors into logical errors such as in the planar Floquet codes constructed in [6], leading to constant distance. Although it is beyond the scope of this current work to develop an algorithm that fully encompasses notions of code distance in Floquet or dynamical codes, it may be fruitful to consider the space-time connectedness of low weight errors for such a task. Generally speaking, we may say that two Pauli operators $\sigma_{\alpha}$ and $\sigma_{\beta}$ are space-time connected in a dynamical code if their difference is a chain of instantaneous stabilizer elements, i.e. $\sigma_{\alpha}=s_{r+t}\dots s_{r}\sigma_{\beta}$. It may be the case that a low weight Pauli error $\sigma_{\beta}$ traverses, in its spacetime path, an undetectable logical error, before evolving into a partially detectable error, whose correction does not correct the previous logical error, or even gets absorbed into the ISG. Such a logical error occurs in [6], for instance. It may therefore be desirable to keep the number of time slices between which stabilizers are measured and used as syndromes as minimal as possible when designing Floquet codes. This intuition partially motivated the distance $\sim d/2$ decoding strategy we described above.
@@
We also note that like the parent Bacon-Shor subsystem code, the Floquet-Bacon-Shor family of codes introduced here do not possess a threshold. The issue is that the weight of the stabilizers used in both examples of codes grows as the distance increases, so that erroneous stabilizer measurements get likelier with increasing distance for any fixed value of independent error probability.
@@ IV.2 Self-correcting errors
Suppose that a 2-qubit $XX$ error occurs on an edge somewhere above cell A between rounds 0 and 1 mod 4. This will flip the two transient, but not the two permanent, $Z$-type stabilizers that straddle this edge, lie along column AD, and are measured in round 1 mod 4. In the very next round 2 mod 4, we carry out a measurement of an $XX$ check operator on precisely the same edge, and the error becomes part of the ISG. In other words, the appropriate correction operator automatically applies itself because of the measurement schedule itself. This occurs even if this 2-qubit error anti-commutes with a logical operator at round 0 mod 4, such as if it occurred on the top edge of cell A. We show this more explicitly in the appendix. Such self-correcting errors may be understood as a residual gauge degree of freedom in the Floquet version of the parent subsystem code, as even though such errors do not belong in the ISG and will produce syndromes, they are nevertheless gauge checks of the parent subsystem code.
@@ IV.3 Multiple Dynamical Qubits
The discussion above generalizes fairly straightforwardly to the case of adding multiple dynamical logical qubits. In this case, the main difference is that there are now more temporary stabilizers used to form detectors in each plaquette row and column. In particular, there is one such detector between any neighboring pair of gauge defects, as well as a detector between any boundary gauge defect and the lattice boundary as before. An even number of $X$ errors between gauge defects in the same plaquette row are now part of the ISG at round 0, and no longer count as errors. Only single qubit $X$ errors between any consecutive pair of gauge defects, or a gauge defect and a lattice boundary count as errors, and similarly for $Z$ errors. It was already noted in Section III.2 that the minimum distance of all ISGs shrinks to roughly $d/\sqrt{k}$ with the addition of $k$ dynamical qubits. The arguments for the single logical qubit discussed above for the code distance generalize to the case of multiple logical qubits as well, and therefore it continues to be the case that the smallest weight of any error operator leading to a trivial syndrome is roughly $d/\sqrt{k}$.
@@ V Conclusions
In this paper, we have described the introduction of dynamical logical qubits to the Bacon-Shor subsystem code by modifying the measurement schedule of the weight 2 check operators that generate the gauge group of the parent sub-system code. Such a measurement schedule introduces and maintains gauge defects throughout the measurement cycle. This broad prescription of identifying the virtual qubit degrees of freedom in a subsystem code, and picking a measurement schedule that can maintain gauge defects across measurement cycles as a way to introduce dynamical logical qubits may be more broadly applicable in constructing more examples of Floquet codes.
We note that an alternative measurement schedule such as the one discussed in this paper can significantly increase the number of logical qubits the code allows for. While the parent Bacon-Shor subsytem code hosts a single logical qubit, the Floquet-Bacon-Shor family of codes can allow $k=\Theta(n)$ many logical qubits, with the tradeoff of a having a constant distance $d=\Theta(1)$. This implies that the Floquet version can saturate the subsystem bound $kd=O(n)$ [38], even though the parent subsystem code does not.
We showed that such an alternative measurement schedule also introduces temporary stabilizers that have smaller weight than the permanent stabilizers of the parent Bacon-Shor subsystem code, but that can nevertheless be used to form detectors and perform error correction using standard matching techniques. We reported the code distance as the minimum weight of any error operator that would yield a trivial syndrome, and noted certain obstacles to forming an all-encompassing notion of Floquet code distance, especially in the fault-tolerant setting. We leave a detailed consideration of fault tolerance, together with fault tolerant implementation of logical gates, in this code to future work.
Lastly, we also note that skipping measurements of some check operators in some rounds was also used in [40] to introduce a threshold to the Bacon-Shor code. In our paper, we used such skipped measurements to introduce dynamical logical qubits to the Bacon-Shor code. We leave it as an open problem to investigate whether one could combine both sets of ideas to introduce a threshold to the Bacon-Shor code in the presence of dynamical logical qubits.
@@ VI Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. M.S.A. acknowledges support from USRA NASA Academic Mission Services under contract No. NNA16BD14C with NASA, with this work funded under the NASA-DOE interagency agreement SAA2-403602 governing NASA’s work as part of the SQMS center. The authors wish to thank Jason Saied and Ryan LaRose for useful discussions on Floquet and error correcting codes in general, as well as Nikolas Breuckmann, Christophe Vuillot and Margarita Davydova for helpeful comments during the final stages of this work. M.S.A. would also like to thank Alex Townsend-Teague and Julio Magdalena de la Fuente for useful discussions, especially involving the ZX calculus, and especially Xiaozhen Fu, for enlightening correspondence on calculating the distance of Floquet codes.
@@ Appendix B Self-correction
Appendix C Code distance
Here, we describe in detail the application of the distance algorithm of [39] to our construction of Floquet-Bacon-Shor codes. In particular, we compute the unmasked distance, defined as the normalizer of certain unmasked stabilizers, excluding a suitably identified gauge group, both of which are identified with the aid of the distance algorithm. The unmasked stabilizers could be thought of as the detectors that can form from the set of elements of a given ISG, while the gauge group roughly corresponds to the set of measurements in the current round and their destabilizers from the next round, both of which would commute with any such detectors, and therefore belong in the normalizer of the unmasked stabilizers, but do not constitute logical errors.
The distance algorithm works by sequentially updating two sets, labelled $V$ and $C$. In addition, we also initialize two empty sets $\tilde{P}$ and $\tilde{U}$. The set $V$ is initialized as the trivial stabilizer group $\langle I\rangle$, and updated according to the rules of ISG growth. In particular, if a Pauli operator $m$ is measured, then
  • (A) If $[m,v]=0$ $\forall v\in V$, then $V$ is updated to $V^{\prime}=\langle V,m\rangle$. Note that $V$ remains unchanged if $m\in V$ already, but increases its rank by 1 if $m\not\in V$.
  • (B) If $\{m,v_{j}\}=0$ for some subset ${v_{j}}\subset V$, then the first of such elements (in any arbitrary ordering) $v_{1}$ is removed from $V$ and the rest updated according to $v_{j}\rightarrow v_{j}v_{1}$, so that $[m,v_{j}v_{1}]=0$. The rank of $V$ remains unchanged in this case.
Meanwhile, the set $C$ is initialized to some given ISG, with respect to which the unmasked distance is computed, and its elements are only ever removed over the course of the algorithm, never added. In particular,
  • (C) If $[m,c]=0$ $\forall c\in C$, then $C$ is left unchanged.
  • (D) If $\{m,c\}=0$, and
    • (i) $[m,v]=0$ $\forall v\in V$, then $c$ is added to the (initially empty) set $\tilde{P}$
    • (ii) $\{m,v\}=0$ for some $v\in V$, then $c$ is updated to $c\rightarrow c\cdot v$ (note that this is similar to how a logical operator is updated)
Finally, at any stage of the measurement schedule, we update the (initially empty) set $\tilde{U}$ that contains two types of elements
  • All elements in $\langle C\rangle\cup\langle V\rangle$,
  • All elements $c\in C$ that get updated to the identity, i.e. $c\rightarrow c\cdot v=\mathbb{I}$ according to rule (D)(ii) above.
Roughly speaking, these correspond to the elements in the original ISG, which was identified with $C$ at the very beginning of the algorithm, that are able to form detectors. These detectors are formed by comparison of the measured values of the same operator in two different rounds, which allows us to readily identify $\tilde{U}=U$, the set of all unmasked stabilizers. Algorithm 2 of [39] also allows for a more general case where detectors can form from measurements across several rounds, and we refer the reader to that paper for further details.
Similarly, the set $\tilde{P}$ contains elements, which along with suitably identified destabilizers according to Alogrithm 3 of [39] allows us to identify the gauge group, which in our case is also simplified in our case due to working with the virtual qubit operators directly.
Figure 6: The distance algorithm of [39] applied to a Floquet-Bacon-Shor code with a single dynamical logical qubit. Unlike in previous figures, here we depict the permanent Bacon-Shor stabilizers in addition to the virtual operators of the gauge qubits for convenience.
We apply this algorithm to an example Floquet-Bacon-Shor code on a $7\times 7$ lattice, computing the unmasked distance of the ISG $S_{4k}$ for $k\in\mathbb{Z}$, as depicted in Fig. 6. In the notation there, the empty grid represents the trivial group $\langle I\rangle$, while colored grids represent a group generated by the denoted virtual qubit operators. The set $V_{0}$ is initialized to the empty grid, while the set $C_{0}$ is identified with the ISG $S_{4k}$.
At each of the rounds, we measure all the virtual qubit operators as depicted, including those for the stabilizer qubits of the Bacon-Shor subsystem code. At round 1, this updates the set $V_{1}$ to exactly equal the measured set $M_{1}$ using rule (A) above, while elements from $C$ are removed using rule (D)(i) and dumped into the set $\tilde{P}_{1}$, which remains unchanged (i.e. $\tilde{P}_{j}=\tilde{P}_{1}$ $\forall j$) after this step. The set $\tilde{U}_{1}$ contains the intersection of $C_{1}$ and $V_{1}$ at this stage.
At round 2, the measured operators update the set $V_{2}$ to almost coincide with the steady state ISG $S_{4k+2}$, except we have not yet measured one $X$-type and one $Z$-type permanent Bacon-Shor stabilizers yet. Thet set $C_{2}$ has all the $Z$ (blue) gauge operators along the defect column updated to the identity using rule (D)(ii), and therefore these now persist in $\tilde{U}_{2}$, which in addition also now contains $X$ (red) gauge operators along a defect row since they belong in the intersection of $C_{2}$ and $V_{2}$.
At round 3, these $X$ (red) gauge operators are updated to the identity using rule (D)(ii) and therefore persist in $\tilde{U}_{3}$, which now also includes all the permanent $Z$ (blue) Bacon-Shor stabilizers, since these belong in the intersection of $C_{3}$ and $V_{3}$. The latter is now almost identical with the steady state ISG $S_{4k+3}$, except it does not contain one of the $X$ (red) permanent Bacon-Shor stabilizers, while the set $C_{3}$ now only contains all the permanent Bacon-Shor stabilizers. Any number of measurements beyond this will not change the structure of $C_{j}$, as it will always contain these, and only these, permanent Bacon-Shor stabilizers for all $j\geq 3$.
Finally, at round 4, completing one measurement cycle, the set $V_{4}$ coincides exactly with the steady state ISG $S_{4(k+1)}$, the previously noted $X$ gauge operators along the defect row in $\tilde{U}_{3}$ persist in $\tilde{U}_{4}$, since they now belong in the intersection of $V_{4}$ and $C_{4}$. The colored generators shown in $\tilde{U}_{4}$ precisely generate the group of unmasked stabilizers $U$.
Meanwhile, the gauge group is generated by the generators shown in $\tilde{P}$ along with a suitable choice of destabilizers. These are depicted in the bottom part of Fig. 6. Altogether, the normalizer of the unmasked stabilizers that exclude all elements in this gauge group give us precisely the distance scaling of $O(d)$ on a $d\times d$ square lattice, as argued previously using a decoding graph picture instead.